OpenStudy (anonymous):

the sum of eight times a number, and 6, is equal to seven times the number

OpenStudy (anonymous):

Set up an equation with your information. 8 times a number=8x. Sum of 8x and 6=8x+6. Seven times the number=7x. Set these things equal to each other, like the problem says. 8x+6=7x Solve for x.

OpenStudy (anonymous):

set up this equation 8x+6=7x and then solve for x